OpenStudy (anonymous):

(sqrt3)/2 + (4sqrt3)/3

OpenStudy (anonymous):

is it\[\sqrt{3}/2+4(\sqrt{3})/3\]

OpenStudy (anonymous):

No, the 4 is\[(4\sqrt{?})\]

OpenStudy (anonymous):

\[\sqrt[4]{3}\]you mean

OpenStudy (anonymous):

Nope, just the 4 infront

OpenStudy (anonymous):

using calculator it is 3.1754 (approx)

OpenStudy (hoblos):

√3/2 + 4√3/3 = 3√3/6 + 8√3/6 (common denominator) = 11√3 /6